OPTICAL ZEEMAN SPECTROSCOPY OF THE A^{2}Π/B^{2}Σ^{+} - X^{2}Σ^{+} BAND SYSTEMS OF CALCIUM MONOHYDRIDE, CaH^{a}

\underline{JINHAI CHEN}, JAMIE GENGLER, and TIMOTHY STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; JOHN M. BROWN, Physical and Theoretical Chemistry Laboratory, University of Oxford, South Parks Road, Oxford, United Kingdom OX1 3QZ.

Numerous branch features in the (0,0) band of the B^{2}Σ^{+} - X^{2}Σ^{+} system and in the (0,0) band of the A^{2}Π - X^{2}Σ^{+} system of calcium monohydride, CaH, have been recorded at near the natural line width limit in the presence of a tunable static magnetic field of up to 1200 Gauss. The observed Zeeman induced shifts were successfully modeled using the traditional effective Hamiltonian to account for the Δv=0 interaction and explicit inclusion of the interaction matrix elements for the perturbations between the (v=1) A^{2}Π and (v=0) B^{2}Σ^{+} states. A comparison is made with the analysis of the Zeeman effect of the (0,0) B^{2}Σ^{+} - X^{2}Σ^{+} R_{1} (0.5) branch feature^{b}’^{c}.

aFunded by NSF-Experimental Physical Chemistry

bB. Friedrich, J.D. Weinstein, R. deCarvalho, and J.M. Doyle, J. Chem. Phys. 110, 2376-2383 (1999).

cR. deCarvalho, J.M. Doyle, B. Friedrich. T. Guillet, J. Kim, D. Patterson, and J.D. Weinstein, Eur. Phys. J. D 7, 289-309